OpenStudy (jaylelile):

Please help, I'm lost! I keep getting 1,5566.20 but that's incorrect. Problem below:

OpenStudy (jaylelile):

\[3x+5y ^{2}+8x ^{2}; x=1.208, y=7.666\]

OpenStudy (anonymous):

just substitute the known values into the function? use a calc?

OpenStudy (jaylelile):

I keep getting the same answer that isn't right though...

OpenStudy (anonymous):

\[3(1.208)+5(7.666)^2+8(1.208)^2\]

OpenStudy (mathstudent55):

\(\large 3\color{red}{x}+5\color{green}{y} ^{2}+8\color{red}x ^{2}; \color{red}{x=1.208}, \color{green}{y=7.666}\) \(\large 3(\color{red}{1.208})+ 5(\color{green}{7.666} ^{2})+8(\color{red}{1.208} ^{2}) \)

OpenStudy (mathstudent55):

Do the squares first.

OpenStudy (mathstudent55):

\(\large 3(\color{red}{1.208})+ 5(\color{green}{7.666} ^{2})+8(\color{red}{1.208} ^{2})\) What do you get for \(7.666^2\)? What do you get for \(1.208^2\)?

OpenStudy (jaylelile):

ohhh I wasn't doing it in the right order.... I got 58.76 and 1.459

OpenStudy (mathstudent55):

\(\large 3(\color{red}{1.208})+ 5(\color{green}{58.767556})+8(\color{red}{1.459264})\)

OpenStudy (mathstudent55):

Now do the 3 multiplications.

OpenStudy (jaylelile):

Do the multiplications and then add. I got 309.17 for my final answer.

OpenStudy (jaylelile):

thank you so much!

OpenStudy (mathstudent55):

\(\large 3(\color{red}{1.208})+ 5(\color{green}{58.767556})+8(\color{red}{1.459264})\) \(= 3.624 + 293.83778 + 11.674112\) \(=309.135892\)

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

I got 99.612 one time, 309.135892 the second.